OpenStudy (anonymous):

Simplify. Express the product as a radical expression. (3rd root of x^2)(4th root of x)

OpenStudy (anonymous):

To make my question clearer.... \[\sqrt[3]{x^2} \times \sqrt[4]{x}\]

OpenStudy (anonymous):

x^(2/3) * x^(1/4) = x^( (2/3) + (1/4) ) = x ^ (11/12) = \[\sqrt[12]{11}\]

OpenStudy (anonymous):

sorry it is gonna be \[\sqrt[12]{x^{11}}\]